OpenStudy (blackstreet23):

How do you evaluate the line integral F*Dr along a line segment C from P (-2,4) to Q (-2,11) if F(X,y) = 2xj?

OpenStudy (zale101):

Do you remember how to paramterize a line? Ex: \(x=at+x_0\) \(y=bt+y_0\)

OpenStudy (zale101):

First step: Paramterize a line and create r(t) equation Second step: Plug the equation you got for x and y in to F(x,y) changing it to F(t) Third: Take the derivative of r(t), and get dr. Fourth: See which t numbers will create P (-2,4) to Q (-2,11) in the r(t) equation. Let us begin with the first step.

OpenStudy (zale101):

To parameterize, you first need to look at the given two points P (-2,4) to Q (-2,11). \(x=at+x_0\) a is the final minus initial. P is the initial point, Q is the final point

OpenStudy (zale101):

so \(x=(final ~x- Initital x)t+x_0\)\)

OpenStudy (zale101):

x0 is and y0 can be replaced by either point P or Q

OpenStudy (zale101):

Paramterization is introduced from the last chapter of calculus 2 or the beginning chapter of calculus 3.

OpenStudy (zale101):

x0 and y0 can be replaced by either point P or Q**

OpenStudy (zale101):

@blackstreet23 does it make sense?

OpenStudy (zale101):

I'll parameteriz x We have two given point Initial =P(-2,4) and Final= Q(-2,11) The x final is -2 the x initial is -2. So a= (-2-(-2)) x0 can be either one of the x values of the given points P or Q. I will choose the x value of P which is -2 for my x0 but you will have to do the same for your y0. You have to choose point P which will have a y0 of 4. So, my x equation is: \(x=(-2(-2))t+(-2)\) \(x=(-2+2)t-2\) \(x=(0)t-2\) \(x=-2\)<-- x equation

OpenStudy (zale101):

Do the same for the equation for y.

OpenStudy (zale101):

and then we are done with the first step.

OpenStudy (blackstreet23):

What is f(t) and r(t) ?

OpenStudy (blackstreet23):

Are you coming tomorrow I am going to sleep ?

OpenStudy (irishboy123):

-28

OpenStudy (irishboy123):

ie -4 * 7